Sir Lancelot found four consecutive even integers such that 10 times the sum of the first and the fourth was 24 greater than 9 times the sum of the second and fourth. what were his integers?
Answer by chessace(471)

10 * (s + s+6) = 24 + 9 * (s+2 + s+6)
20 * s + 60 = 24 + 18 * s + 9 * 8
2 * s = 24 + 72 - 60 = 36
s = 18
Numbers are 18 20 22 24